OpenStudy (anonymous):

What is the solution set for 2x^2 +14x = 35

OpenStudy (anonymous):

x =(-14-√476)/4=(-7-√ 119 )/2= -8.954 x =(-14+√476)/4=(-7+√ 119 )/2= 1.954